OpenStudy (anonymous):

Express the volume of a cone, V, as a function of its radius, r, if the radius is 1/3 of the height. I have tried (1/9)*pi*r^2 but it did not work.

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 } *\pi r^2h\] r=h/3 h=3r plug in this equation for h and you will get your answer

OpenStudy (anonymous):

So would the answer be V=pi^3?

OpenStudy (anonymous):

that's what i got

OpenStudy (anonymous):

oops V=pi*r^3

OpenStudy (anonymous):

Thanks! Haha yeah I forgot to put in the r oops